OpenStudy (anonymous):

(2square root7a)(-4square root7a) solve

OpenStudy (anonymous):

2sqrt(7a)*(-4sqrt(7a)) = 2*sqrt(7a)*(-4)*(sqrt(7a)) , which is the associative property of multiplication. = sqrt(7a)*sqrt(7a)*(2*(-4)) ,which is the commutative property of multiplication Do you know how to solve that?

OpenStudy (anonymous):

I am using hte foil method but Iam stuck when I get multiplying the 2 square roots of 7a together---2(-4) + 2sqrt7a-4sgrt7a +sgrt7asquared does this make since what I am doing

OpenStudy (anonymous):

You can't use foil for this. foil is only for (a+b)(c+d). Not for (a*b)*(c*d)

OpenStudy (anonymous):

so is my answer -8sqrt49asquared

OpenStudy (anonymous):

(a*b)*(c*d) = a*b*c*d = abcd = acbd = bdca It's both the associative property and commutative property combined

OpenStudy (anonymous):

Wrong, but close

OpenStudy (anonymous):

See, sqrt(7a)*sqrt(7a) = sqrt(7a*7a) = sqrt( (7a)^2 )

OpenStudy (anonymous):

-8sqrt7a^2

OpenStudy (anonymous):

what is sqrt(x^2) ?

OpenStudy (anonymous):

\[2(-4) +2\sqrt{7a}-4\sqrt{7a}+\sqrt{7a ^{2}}\] I get this far and then I am stuck

OpenStudy (anonymous):

That is wrong. You can't use foil. This is pure multiplication, unless you stated your original problem wrong.

OpenStudy (anonymous):

the question reads perform the indicated operations and simplify your answer. write your answer in simple radical form\[(2\sqrt{7a})(-4\sqrt{7a)}\]

OpenStudy (anonymous):

Yes but this is multiplication

OpenStudy (anonymous):

so you can't use foil

OpenStudy (anonymous):

-8\[-8\sqrt{49a ^{2}}\] equals 8*7a=56a my answer is 56a

OpenStudy (anonymous):

oops -56a

OpenStudy (anonymous):

Which is 99% correct

OpenStudy (anonymous):

It is probably good enough for your teacher, but a true mathematician would say this is wrong.

OpenStudy (anonymous):

the answer is -56a if a is positive, and 56a if a is negative, because the square root sign is actually a principal square root sign, and it must be positive.